Near the surface of Earth, an electric field points radially downward and has a magnitude of approximately 100 N/C. 1) What charge in mC (magnitude and sign) would have to be placed on a penny that has a mass of 3.11 g to cause it to rise into the air with an upward acceleration of 1.990 m/s2? (Express your answer to three significant figures.)
Answer #1

1 - 9E - mg = mxa. э - TE = m(t) з – бих мою 3. 1 x 102, ( 1 - 110 + 1- 3 ) 10-4 = y - C - 3.61x [o 1 Answer